### v1

# ProRefine: Inference-time Prompt Refinement with Textual Feedback

Abstract
Agentic workflows, where multiple AI agents collaborate to accomplish complex tasks like reasoning or planning, are becoming increasingly prevalent. However, these workflows often suffer from error propagation and sub-optimal performance, largely due to poorly designed prompts that fail to effectively guide individual agents. This is a critical problem because it limits the reliability and scalability of these powerful systems. We introduce ProRefine, an innovative inference-time prompt optimization method that leverages textual feedback from large language models (LLMs) to address this challenge. ProRefine dynamically refines prompts for multi-step reasoning tasks without additional training or ground truth labels. Evaluated on five benchmark mathematical reasoning datasets, ProRefine significantly surpasses zero-shot Chain-of-Thought baselines by 3 to 37 percentage points. This approach not only boosts accuracy but also allows smaller models to match the performance of larger ones, highlighting its potential for efficient and scalable AI deployment, and democratizing access to high-performing AI.
1 Introduction
A critical condition for the advancement of artificial intelligence (AI) is that AI outputs align with human values [10]. Human feedback, a scarce and subjective resource [11, 14, 26, 44], plays a vital role. LLMs, trained on vast datasets, utilize alignment techniques to generate more human-like and accurate responses [16, 30, 37]. Traditional methods employ Reinforcement Learning from Human Feedback (RLHF) [5], with Proximal Policy Optimization (PPO) [33] being a popular choice for fine-tuning LLMs.
Recent research has explored using LLMs as judges and evaluators [3, 4, 17, 20, 43, 49]. Building on use of LLMs to judge outputs, are compound systems that use LLMs to generate feedback / critiques [27, 31]. Akyurek et al. [1] and Wadhwa et al. [42] explore agentic frameworks to detect, critique, and refine tasks that require factual correctness. Another approach is to iteratively refine output using self-generated feedback [21]. TextGrad [48] introduced automatic “differentiation” via text, using textual feedback to optimize the performance of individual components within a compound AI system.
Figure 1: Overview of the ProRefine system, illustrating the iterative process of prompt optimization using feedback from LLMs in an agentic workflow.
Our work focuses on optimizing the prompt, a key element in chain-of-thought (CoT) [45] based LLM reasoning. CoT mimics human problem-solving by breaking down complex tasks into smaller, manageable steps. The system we introduce here, ProRefine (Inference-time Prompt Refinement with Textual Feedback), builds upon CoT by adaptively improving prompts using feedback ($LLM_{feedback}$) and an optimizer ($LLM_{optimizer}$) to refine prompts for the task-performing LLM ($LLM_{task}$). This workflow (Figure 1), motivated by the teacher-student framework [39] where a teacher agent guides a student agent to perform a task by providing feedback at intermediate steps, but implemented via LLM interactions without pre-training, represents a novel approach to adaptive agentic reasoning. We explore policy optimization for aligning compound AI systems, drawing inspiration from TextGrad and policy gradient algorithms like PPO. This work investigates the following research questions.
RQ1
Can textual feedback be used to improve the performance of LLMs at inference-time?
RQ2
Does model size impact the ability of LLMs to utilize textual feedback?
RQ3
What is the impact of using a verifier at inference-time?
While prior work has explored prompt optimization [7, 8, 34, 46], they all require training examples to produce task-specific prompts. By contrast, ProRefine is task agnostic and requires no additional training or ground-truth labels. ProRefine is an inference-time optimization method that relies on test-time compute availability and the ability of LLMs to provide and act upon feedback for optimization, see Figure 2 for an example. This offers a valuable alternative to solely relying on scaling up the base model size for all queries, especially in scenarios where the smaller model suffices for most tasks but needs enhancement for specific critical ones.
ProRefine’s ability to break complex tasks into smaller steps and self-improve using feedback makes it a practical solution for multi-step agentic workflows and resource-constrained environments where deploying the largest models isn’t feasible, but temporary access to a capable feedback LLM (perhaps via a separate API call) is possible for critical tasks. This method is also suitable for black-box LLMs, where the internal parameters are inaccessible and only API access is available. ProRefine could prove to be crucial in situations demanding higher interpretability, where the feedback steps (outputs of $LLM_{feedback}$) offer insights into the reasoning correction process, and applications requiring dynamic adaptation without retraining/fine-tuning cycles. To demonstrate the effectiveness of ProRefine, we evaluate our method across five benchmark mathematical reasoning datasets.
Key Contributions:
  • We propose a novel method - ProRefine for inference-time prompt optimization using textual feedback.
  • We evaluate ProRefine on five datasets: object counting, word sorting, grade-school math problem-solving, math word problems, and algebraic word problems, and compare our method against CoT and TextGrad.
  • We evaluate the importance of using a verifier at inference-time.
Figure 2: ProRefine example. Given an input query (which in this case has correct answer: 13), and an initial prompt, the task model ($LLM_{task}$) gives an incorrect answer. Prorefine uses two additional models, $LLM_{feedback}$ and $LLM_{optimizer}$, to iteratively improve the prompt as the $LLM_{task}$ generates its response. Refining the prompt during generation allows the feedback model to target local regions of the response, providing finer-grained feedback than waiting for the response to complete.
2 Related Work
ProRefine draws inspiration from and contributes to several interconnected research areas. The performance of LLMs is heavily dependent on the quality of the prompts they receive. Early efforts in this domain centered on manual prompt crafting [45], a meticulous process of designing effective prompts to elicit desired responses. Recognizing the limitations and scalability challenges of manual methods, research has increasingly focused on automatic prompt optimization.
Some pioneering automatic methods, such as AutoPrompt [34] and RLPrompt [7] , employ gradient-based search and reinforcement learning techniques, respectively. AutoPrompt [34] for instance, uses gradient-based search to generate prompts for masked language models. It reformulates tasks as fill-in-the-blank problems, achieving performance comparable to supervised models in tasks like sentiment analysis. However, it requires training data and gradient access, limiting its applicability to black-box models. Other approaches leverage LLMs themselves for prompt generation [22, 27, 46, 47, 50]. ProRefine distinguishes itself by operating at inference-time without training data or ground truth labels, extending the concept of Chain-of-Thought prompting [45] by iteratively refining the prompt.
There is a substantial and growing body of work exploring the capacity of LLMs to act as judges or evaluators [3, 4, 17, 20, 41, 43, 49, 51] of text quality, correctness, and other attributes. This capability is crucial for automating various aspects of model development and deployment, from assessing generation quality to providing feedback for model improvement. ProRefine utilizes this concept, employing an LLM to provide feedback (a form of evaluation), but crucially uses this feedback for optimization and combines in a novel fashion inference-time operations with LLM-generated feedback.
The idea of LLM self-improvement and iterative refinement is also highly relevant. Self-Refine [21] is a prominent example, where an LLM generates both output and feedback, using the latter for refinement. Other works explore self-critiquing [31] and reinforcement learning for critique generation (RL4F) [1], along with various feedback and refinement mechanisms [8, 15, 28, 29, 32, 42]. ProRefine shares this spirit but focuses specifically on prompt refinement, making it suitable for agentic workflows and black-box LLMs, unlike direct output modification.
ProRefine builds upon the core ideas in TextGrad [48], which uses textual feedback for optimization within compound AI systems. However, ProRefine applies feedback at intermediate steps for inference-time prompt optimization and rapidly evolving tasks where fine-tuning is infeasible, a key distinction. Moreover, TextGrad is a supervised fine-tuning method, while ProRefine requires no training data, making its data requirements less onerous. Finally, ProRefine belongs to the broader category of inference-time methods [24, 35], that improve LLMs without weight modification, such as multi-agent debate [9]. Inference-time methods aim to improve the performance of models by utilizing test-time compute resources. Recent advances in understanding reasoning in LLMs [12, 13] also provide context to our prompt improvement.
In essence, ProRefine uniquely integrates inference-time operation, LLM-generated feedback, and prompt optimization to enhance multi-step reasoning in agentic workflows, effectively addressing limitations of prior work related to reliance on training data and applicability to black-box models.
3 ProRefine
ProRefine is an inference-time prompt optimization algorithm that optimizes prompts by using textual feedback. ProRefine involves interactions between three LLMs:
$LLM_{task}$: Executes the task based on the current prompt, generating the initial and subsequent outputs.
$LLM_{feedback}$: A model that critiques the $LLM_{task}$’s output, providing detailed feedback on improvements. This model should be capable of providing insightful and accurate critiques [2, 31].
$LLM_{optimizer}$: Interprets the feedback and refines the prompt, aiming for coherent and task-focused improvements. This LLM is crucial for ensuring the prompt evolves effectively.
ProRefine
(Algorithm 1) works as follows:
Input: Query: $q$, Initial prompt: $p$, tokens_per_step: $k$, max_steps: $n$, LLMs: $LLM_{task}$, $LLM_{feedback}$, $LLM_{optimizer}$
Output: Optimized prompt: $p^{*}$
$p^{*}=p$
for $i=1$ to $n$ do
       $o_{i}=LLM_{task}(p^{*},q)$ // Generate $i*k$ tokens
       $f_{i}=LLM_{feedback}(q,o_{i})$ // Get textual feedback
       $p*=LLM_{optimizer}(p^{*},f_{i})$ // Optimize the prompt
       if $EOS\_token$ in $o_{i}$ then
            break
return $p^{*}$ // Return final optimized prompt
Algorithm 1 ProRefine
Initialization:
Start with an initial prompt $p$ for the task, a query $q$, and parameters defining the generation and optimization process ($k$ tokens per step, $n$ maximum steps).
Generation and Feedback Loop:
  • Generation: Use $LLM_{task}$ to generate an output based on the current prompt $p^{*}$ and query $q$. This step is limited to $k$ tokens to control the granularity of feedback.
  • Feedback: $LLM_{feedback}$ evaluates the generated output $o_{i}$ against the query $q$ to provide textual feedback $f_{i}$. This feedback encapsulates how the output could be improved, focusing on aspects such as accuracy, relevance, or coherence.
  • Optimization: $LLM_{optimizer}$ uses the feedback $f_{i}$ to refine the prompt $p^{*}$. This step involves modifying the prompt to better align with the task requirements or to correct identified deficiencies in previous generations.
Termination:
The process iterates until either the maximum number of steps $n$ is reached or an end-of-sequence (EOS) token is detected in the output, indicating the completion of the task.
3.1 Unifying Verifier and Feedback
At inference time, verifiers play a crucial role in judging model outputs [6, 18, 36]. Often trained on extensive corpora, these verifiers can be outcome-based, assessing the final result’s correctness, or process-based, evaluating the validity of intermediate steps [18, 36, 40]. For simplicity in this study, we do not train a bespoke verifier; rather, we employ the Llama3.1-70B-instruct model to function as both the feedback mechanism ($LLM_{feedback}$) and the verifier. A smaller model, specifically fine-tuned for these tasks, could also be used. The verifier’s function is to evaluate the initial output generated by $LLM_{task}$ for each query. If this output is assessed as incorrect, the refinement process is triggered; otherwise, the output is used directly.
To quantify the verifier’s impact, we analyze three distinct scenarios: ProRefine (verifier), our standard approach which employs $LLM_{feedback}$ to guide refinement; ProRefine (no verifier), wherein the refinement process operates without verifier input; and ProRefine (optimal verifier), guided by a perfect verifier (simulated using ground-truth labels). This optimal condition reveals the upper bound of the refinement loop’s potential. Consequently, the performance difference between ProRefine (verifier) and ProRefine (optimal verifier) underscores the significance of verifier accuracy. It is important to note that ProRefine’s methodology does not inherently rely on labels or optimal verification, despite their use in this specific evaluation.
3.2 Implementation Details
The granularity and duration of the optimization process are governed by two parameters: $k$, the number of tokens per step, and $n$, the maximum number of steps. These parameters can be adjusted based on the complexity of the task and the desired output quality. For example, rather than generating feedback every $k$ tokens, we might instead choose to provide feedback after each sentence or paragraph, particularly in tasks such as machine translation or text summarization, where larger semantic units may be more meaningful. We run all of our experiments on compute nodes with a single Nvidia A100 GPU (80GB), 24 core processor, and 220GB RAM. For model’s generate function we set the following parameters: num_return_sequences=1, do_sample=False, top_p=None, temperature=None.
4 Experiments and Evaluation
4.1 Data
We evaluate ProRefine on five reasoning tasks, each of which involves multi-step reasoning, making them suitable for evaluating prompt optimization in agentic workflows: object counting and word sorting from the BIG-Bench Hard benchmark [38], grade-school math problem-solving from GSM8K [6], math word problems from SVAMP [25], and algebraic word problems from AQUARAT [19]. We use the same dataset splits and evaluation as Yuksekgonul et al. [48] for object counting, word sorting, and GSM8K. Refer to Table 1 for dataset statistics.
Dataset Train Validation Test
Object Counting 50 100 100
Word Sorting 50 100 100
GSM8K 200 300 1319
SVAMP 2516 622 1000
AQUARAT 97467 254 254
Table 1: Datasets used for evaluating ProRefine on multi-step reasoning tasks.
Dataset Method Llama-3.2 1B-it Llama-3.2 3B-it Llama-3.1 8B-it
Object Counting CoT 0.48 [0.382, 0.578] 0.65 [0.556, 0.744] 0.73 [0.643, 0.817]
TextGrad 0.62 [0.524, 0.716] 0.73 [0.643, 0.817] 0.86 [0.792, 0.928]
ProRefine (no verifier) 0.51 [0.412, 0.608] 0.75 [0.665, 0.835] 0.77 [0.687, 0.853]
ProRefine (verifier) 0.6 [0.503, 0.696] 0.72 [0.632, 0.808] 0.89* [0.839, 0.959]
ProRefine (optimal verifier) 0.67 [0.577, 0.763] 0.85* [0.780, 0.920] 0.94* [0.893, 0.987]
Word Sorting CoT 0.11 [0.048, 0.172] 0.10 [0.041, 0.159] 0.50 [0.401, 0.598]
TextGrad 0.33* [0.237, 0.423] 0.61* [0.514, 0.706] 0.69* [0.599, 0.781]
ProRefine (no verifier) 0.22 [0.138, 0.302] 0.47* [0.372, 0.568] 0.68 [0.595, 0.779]
ProRefine (verifier) 0.19 [0.113, 0.267] 0.32* [0.228, 0.412] 0.71* [0.621, 0.799]
ProRefine (optimal verifier) 0.29* [0.192, 0.368] 0.53* [0.432, 0.628] 0.86** [0.792, 0.928]
GSM8K CoT 0.450 [0.423, 0.476] 0.809 [0.787, 0.829] 0.819 [0.797, 0.839]
TextGrad 0.463 [0.436, 0.489] 0.801 [0.779, 0.822] 0.864* [0.845, 0.882]
ProRefine (no verifier) 0.636** [0.610, 0.662] 0.797 [0.774, 0.818] 0.843 [0.823, 0.863
ProRefine (verifier) 0.654** [0.627, 0.678] 0.866** [0.847, 0.883] 0.885* [0.868, 0.902]
ProRefine (optimal verifier) 0.725** [0.701, 0.749] 0.904** [0.888, 0.920] 0.936** [0.922, 0.949]
SVAMP CoT 0.689 [0.66, 0.718] 0.869 [0.848, 0.890] 0.854 [0.832 , 0.876]
TextGrad 0.684 [0.655, 0.713] 0.861 [0.840, 0.882] 0.84 [0.817, 0.863]
ProRefine (no verifier) 0.774** [0.748, 0.800] 0.878 [0.858, 0.898] 0.877 [0.857, 0.897]
ProRefine (verifier) 0.808** [0.784, 0.832] 0.896 [0.877, 0.915] 0.893* [0.874, 0.912]
ProRefine (optimal verifier) 0.861** [0.840, 0.882] 0.925** [0.909, 0.941] 0.938** [0.923, 0.953]
AQUARAT CoT 0.259 [0.202, 0.31] 0.563 [0.498, 0.620] 0.586 [0.522, 0.643]
TextGrad 0.311 [0.250, 0.364] 0.524 [0.462 , 0.585] 0.559 [0.494, 0.616]
ProRefine (no verifier) 0.205 [0.151, 0.250] 0.343 [0.284, 0.401] 0.398 [0.337 , 0.458]
ProRefine (verifier) 0.268 [0.209, 0.318] 0.551 [0.486 , 0.608] 0.606 [0.542, 0.663]
ProRefine (optimal verifier) 0.354 [0.292, 0.409] 0.598 [0.538, 0.659] 0.657 [0.595, 0.712 ]
Table 2: Test Accuracy with 95% confidence intervals across five benchmark datasets and models. * and ** denote statistically significant improvements over one or two baseline methods, respectively. Results in bold indicate the highest accuracy for a dataset-method combination. demonstrates the upper bound potential of the optimization loop and the impact of verifier quality. Llama3.1-70B-instruct is employed for feedback generation, prompt optimization, and evaluation.
Figure 3: Test Accuracy [with 95% confidence interval] across different models and datasets. Llama3.1-70B-instruct is employed for feedback generation, prompt optimization, and evaluation.
4.2 Experimental Setup
We experiment with three models - Llama3.2-1B-instruct, Llama3.2-3B-instruct, and Llama3.1-8B-instruct [23] for $LLM_{task}$. The prompts are optimized using Algorithm 1, with Llama3.1-70B-instruct used for feedback generation, prompt optimization, and evaluation. We select the values of hyperparameters $k=10$ and $n=25$ to control the granularity of feedback and duration of optimization. Hyperparameters $k$ and $n$ were fixed based on general preliminary exploration and not tuned per task using benchmark training/validation data.
We compare the performance of our method against the zero-shot Chain-of-Thought (CoT) baseline and TextGrad [48], and report test accuracy with 95% confidence interval. We choose TextGrad as a baseline because Yuksekgonul et al. [48] reported performance at par or better than DSPy [15] for prompt optimization on object counting, word sorting, and GSM8k datasets. It is essential to remember that TextGrad is a supervised fine-tuning method that utilizes both the training and validation sets. For TextGrad, we use a comparative setup consisting of a task model to be fine-tuned and Llama3.1-70B-instruct model for feedback generation and backpropagation. The results are shown in Table 2 and Figure 3.
4.3 Results
Our results (Table 2) demonstrate that ProRefine significantly improves $LLM_{task}$ performance over the zero-shot CoT baseline in all but one experiment, and it outperforms TextGrad in 11 out of 15 cases overall. For Llama3.2-1B-instruct model, ProRefine can significantly outperform CoT and TextGrad on 2 out of 5 datasets. For Llama3.2-3B-instruct model, ProRefine can outperform CoT and TextGrad on 3 out of 5 datasets with one significant result. For Llama3.1-8B-instruct model, ProRefine can outperform CoT and TextGrad on all 5 datasets with 4 significant results.
Object Counting
ProRefine improves performance by $3-16$ percentage points over CoT, with significant gains observed for Llama3.1-8B-instruct. It outperforms TextGrad on 2 out of 3 models, yielding a $2-3$ percentage point advantage. However, a performance drop of $2$ points is observed for Llama3.2-1B-instruct.
Word Sorting
Performance gains over CoT range from $8-37$ percentage points, with significant improvements for Llama3.2-3B-instruct and Llama3.1-8B-instruct. ProRefine surpasses TextGrad on 1 of 3 models with a $2$-point gain, but performance drops of $11-14$ points are observed for Llama3.2-1B-instruct and Llama3.2-3B-instruct.
GSM8K
ProRefine achieves $2.4-20.4$ percentage points improvement over CoT, with significant improvement observed for all the models; however, a slight performance drop ($1.2$) is observed for Llama3.2-3B-instruct. It outperforms TextGrad on all models, achieving a $2.1-19.1$ percentage point gain with significant results observed for Llama3.2-1B-instruct and Llama3.2-3B-instruct models. Minor performance drop of $0.4-2.1$ is observed for Llama3.2-3B-instruct and Llama3.1-8B-instruct.
SVAMP
Performance improves by $0.9-11.9$ percentage points over CoT, with significant gains for Llama3.2-1B-instruct and Llama3.1-8B-instruct. ProRefine outperforms TextGrad across all models, with $1.7-12.4$ percentage point gains and significant results for Llama3.2-1B-instruct.
AQUARAT
Gains over CoT range from $0.9-2$ percentage points, but declines of $5.4-22$ points are also observed. ProRefine exceeds TextGrad on 2 of 3 models, with $2.7-4.7$ percentage point gains, though performance drops of $10.6-18.1$ points are also recorded.
Our results demonstrate that using ProRefine with an optimal verifier significantly improves performance across all tasks, achieving the best results in 13 out of 15 cases, highlighting the critical role of verifier quality. Notably, the number of significant improvements increases with larger model sizes.
We also observe that ProRefine can improve the performance of lesser capable models Llama3.2-3B-instruct and Llama3.1-8B-instruct to that of more capable ones Llama3.1-8B-instruct and Llama3.1-70B-instruct respectively. In our experiments using Llama3.1-70B-instruct for $LLM_{task}$, some experiments yielded an accuracy of 1, suggesting potential overfitting. Additionally, we experimented using ProRefine with a comparatively smaller model (Llama3.1-8B-instruct) for $LLM_{feedback}$ and $LLM_{optimizer}$. However, this configuration did not consistently yield substantial improvements and, in some cases, even degraded performance. This aligns with prior findings indicating that smaller models are less capable of providing high-quality feedback [31, 2].
Figure 4: Average number of prompt refinement iterations.
5 Discussion
Regarding RQ1, the results demonstrate that ProRefine is a broadly applicable method utilizing textual feedback for improving LLM performance at inference-time. The “performance gap bridging” effect is particularly noteworthy, suggesting that ProRefine can be an effective alternative to simply scaling up model size, potentially avoiding costly fine-tuning, which could prove critical in resource-constrained settings.
The largest performance gains are observed on the word sorting task, indicating that tasks requiring more complex reasoning or manipulation of intermediate outputs benefit the most from ProRefine’s iterative refinement. The mixed results when using a smaller model for $LLM_{feedback}$ illustrate the importance of “knowledge asymmetry,” i.e., that the feedback model should be “sufficiently capable” of providing useful critiques.
Regarding RQ2, the results indicate that ProRefine outperforms the baselines on 2 and 3 datasets when using the Llama3.2-1B-instruct and Llama3.2-3B-instruct models, respectively, and on all 5 datasets when using the Llama3.1-8B-instruct model. This suggests that performance improvements scale with model size. These findings imply that larger models are preferable to smaller ones, particularly in agentic workflows that may require test-time scaling and the effective use of textual feedback to solve complex tasks.
Regarding RQ3, the results highlight that employing a high-quality verifier is crucial for significantly improving task performance at inference time. Beyond enhancing performance, the verifier also reduces computational cost during inference by guiding the refinement process. Moreover, it opens up promising avenues for future work, where an optimizer could be designed to maximize rewards guided by the verifier’s assessments.
ProRefine can offer a degree of interpretability by exposing the outputs from $LLM_{feedback}$, allowing insights into the model’s reasoning process. Figures 2 and 5 demonstrate cases where model feedback successfully improves the output and where it fails, respectively. Although evaluated on reasoning and math tasks, ProRefine is general and applicable to other tasks.
While ProRefine improves accuracy, it does increase the computational cost during inference. Each step requires calls to both the task LLM and the feedback/optimization LLMs. The trade-off between accuracy and computational cost should be considered when deploying ProRefine for test-time scaling. However, the cost is still considerably lower than full model retraining or fine-tuning. Figure 4 shows the average number of prompt refinement iterations using our standard approach, ProRefine (Verifier), for all datasets and models. Even though we use the same $n=25$ for all datasets and models, Figure 4 suggests that the granularity of the feedback ($k$) and the length of optimization ($n$) may be tuned for specific tasks to achieve an optimal balance between performance and computational cost.
Figure 5: ProRefine example. Given an input query and an initial prompt, the task model ($LLM_{task}$) gives an incorrect answer. Prorefine uses two additional models, $LLM_{feedback}$ and $LLM_{optimizer}$, to iteratively improve the prompt as the $LLM_{task}$ generates its response. Refining the prompt during generation allows the feedback model to target local regions of the response, providing finer-grained feedback than waiting for the response to complete. However, the optimizer is unable to utilize the feedback in this case.
6 Limitations and Future Work
This work has the following limitations that we acknowledge have potential for future explorations:
  • Computational Cost and Scalability: ProRefine’s iterative process increases computational cost compared to single-shot prompting. Scaling to substantially larger or more complex tasks presents both computational and methodological challenges.
  • Generalizability: Evaluation is currently limited to reasoning and mathematical tasks. Further research is needed to assess performance across a broader range of NLP tasks and domains.
  • Model Dependence and Feedback Quality: ProRefine relies on a capable LLM for feedback, which may limit accessibility. The quality and potential biases of this feedback are crucial factors impacting performance and require further investigation.
  • Evaluation and Bias: Using an LLM for evaluation introduces potential biases. More comprehensive human evaluations and robust methods for mitigating evaluator bias are needed.
  • Hyperparameter and Feedback Loop: Current approach requires manual hyperparameter tuning, and iterative prompting might cause issues, and automatic ways need further research.
7 Conclusion
We introduced ProRefine, a novel, practical, and inference-time prompt optimization method for agentic workflows. ProRefine leverages LLM-generated textual feedback to dynamically refine prompts, leading to significant performance improvements without requiring additional training or ground-truth labels. The results across diverse datasets demonstrate ProRefine’s effectiveness for multi-step reasoning tasks and its ability to bridge the performance gap between smaller and larger LLMs. The inference-time nature makes ProRefine readily deployable and adaptable, contributing to more efficient, transparent, and accessible AI systems.